OpenStudy (anonymous):

If f(x) = 1 over 2x + 2 and g(x) = −3x − 1, find g[f(4)]. −19 11 −13 −11

OpenStudy (anonymous):

\[f(x)=\frac{1}{2x+2}\] f(4)=???

OpenStudy (anonymous):

i still do not knw bro

OpenStudy (anonymous):

\[\frac{1}{2(4)+2}=???????\]

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

\[\frac{1}{2(4)+2}=\frac{1}{8+2}=\frac{1}{10}\] how in the world do you get 1/10= 10...

OpenStudy (anonymous):

cuz when u divide 1 and 10 it equls 10

OpenStudy (anonymous):

\[\frac{10}{1}=10\] \[\frac{1}{10}=.1\]

OpenStudy (anonymous):

but thats not an option

OpenStudy (tux):

-13/10